OpenStudy (anonymous):

find the degree measure of a sector in a circle if the radius of the circle is 4 yards and the area of the sector is 4 pi yd ^2

ganeshie8 (ganeshie8):

hint : 360 degrees represent an area of "pi*4^2"

ganeshie8 (ganeshie8):

so u can setup a proportion

ganeshie8 (ganeshie8):

\(\large \frac{360}{\pi *4^2} = \frac{x}{4 \pi}\)

ganeshie8 (ganeshie8):

\(\large \frac{360}{16\pi } = \frac{x}{4 \pi}\)

ganeshie8 (ganeshie8):

you can solve \(x\)